# The Feasibility of Topic-Based Watermarking on Academic Peer Reviews

### v1

@@ Appendix E Class-Specific Classifier Evaluation
(b) BERT Base TBW
(c) RoBERTa Base NW
(d) RoBERTa Base TBW
(e) BERT Few-shot NW
(f) BERT Few-shot TBW
(g) RoBERTa Few-shot NW
(h) RoBERTa Few-shot TBW
(i) BERT Fine-tuned NW
(j) BERT Fine-tuned TBW
(k) RoBERTa Fine-tuned NW
(l) RoBERTa Fine-tuned TBW
Figure 11: Confusion matrices comparing topic-based watermarking (TBW) at $\tau=0.7$ with unwatermarked (NW) text across all model configurations. Each matrix reports the performance of either the BERT or RoBERTa classifier applied to outputs from three LLM variants: base, few-shot, and fine-tuned. Results highlight class-wise prediction behavior across watermarking and classifier settings.
Appendix G Peer Review Shifts Under Paraphrasing
@@
Classifier Model Watermark PEGASUS Shifts DIPPER Shifts
BERT Base NW 58 54
TBW-0.7 37 23
TBW-0.3 51 45
Few-shot NW 24 14
TBW-0.7 24 24
TBW-0.3 24 22
Fine-tuned NW 27 20
TBW-0.7 15 15
TBW-0.3 25 15
RoBERTa Base NW 13 9
TBW-0.7 23 25
TBW-0.3 16 19
Few-shot NW 30 13
TBW-0.7 27 22
TBW-0.3 25 20
Fine-tuned NW 24 14
TBW-0.7 18 22
TBW-0.3 21 18
Table 9: Number of review classification shifts under paraphrasing attacks. Each entry reflects the count (out of 100 paraphrased samples) where the predicted class label differs from the original. Results are grouped by classifier, model variant, and watermarking scheme (NW, TBW-0.7, TBW-0.3), and evaluated separately under PEGASUS and DIPPER paraphrasing models.

### v2

@@ F.4 Class-Specific Classifier Evaluation
Table 10 reports the classification metrics for each classifier and LLM model variant under TBW with $\tau=0.3$. While overall performance remains comparable to the $\tau=0.7$ condition, we observe that the fine-tuned model achieves the highest accuracy across both BERT and RoBERTa classifiers, suggesting that domain adaptation remains a dominant factor in attribution effectiveness even under relaxed topic alignment. This analysis underscores the relative semantic distinctiveness of strongly positive (accept) and moderate (borderline) reviews, while highlighting the challenges involved in distinguishing lower-quality (reject) reviews, which often exhibit more linguistic and structural variability across different watermarking configurations and similarity thresholds.
@@
(b) BERT Few-shot NW
(c) BERT Fine-tuned NW
(d) BERT Base TBW ($\tau=0.7$)
(e) BERT Few-shot TBW ($\tau=0.7$)
(f) BERT Fine-tuned TBW ($\tau=0.7$)
(g) BERT Base TBW ($\tau=0.3$)
(h) BERT Few-shot TBW ($\tau=0.3$)
(i) BERT Fine-tuned TBW ($\tau=0.3$)
(j) RoBERTa Base NW
(k) RoBERTa Few-shot NW
(l) RoBERTa Fine-tuned NW
(m) RoBERTa Base TBW ($\tau=0.7$)
(n) RoBERTa Few-shot TBW ($\tau=0.7$)
(o) RoBERTa Fine-tuned TBW ($\tau=0.7$)
(p) RoBERTa Base TBW ($\tau=0.3$)
(q) RoBERTa Few-shot TBW ($\tau=0.3$)
(r) RoBERTa Fine-tuned TBW ($\tau=0.3$)
Figure 9: Comparison of confusion matrices across different model configurations and watermarking settings. Top row shows BERT results, middle row shows RoBERTa results, with columns representing no watermarking (NW), topic-based watermarking at $\tau=0.7$, and topic-based watermarking at $\tau=0.3$.
@@
F.5 Peer Review Shifts Under Paraphrasing
@@
Classifier Model Watermark PEGASUS DIPPER
Shifts Shifts
BERT Base No watermark 58 54
TBW ($\tau=0.7$) 37 23
TBW ($\tau=0.3$) 51 45
Few-shot No watermark 24 14
TBW ($\tau=0.7$) 24 24
TBW ($\tau=0.3$) 24 22
Fine-tuned No watermark 27 20
TBW ($\tau=0.7$) 15 15
TBW ($\tau=0.3$) 25 15
RoBERTa Base No watermark 13 9
TBW ($\tau=0.7$) 23 25
TBW ($\tau=0.3$) 16 19
Few-shot No watermark 30 13
TBW ($\tau=0.7$) 27 22
TBW ($\tau=0.3$) 25 20
Fine-tuned No watermark 24 14
TBW ($\tau=0.7$) 18 22
TBW ($\tau=0.3$) 21 18
Table 11: Number of review classification shifts under paraphrasing attacks. Each entry reflects the count (out of 100 paraphrased samples) where the predicted class label differs from the original. Results are grouped by classifier, model variant, and watermarking scheme (NW, TBW ($\tau=0.7$), TBW ($\tau=0.3$)), and evaluated separately under PEGASUS and DIPPER paraphrasing models.